OpenStudy (anonymous):

PLEASE HELP!!! Which of the following is equal to square root of the cube root of 5? 5 to the power of 1 over 3 5 to the power of 1 over 6 5 to the power of 2 over 3 5 to the power of 3 over 2

OpenStudy (tkhunny):

Use your form and exponent rules. \(\sqrt{\sqrt[3]{5}} = \left(5^{1/3}\right)^{1/2} = 5^{1/3\cdot 1/2} = 5^{1/6}\)

OpenStudy (anonymous):

Thank you :)

OpenStudy (tkhunny):

Okay, now you do one.